OpenStudy (anonymous):

The primary coil of a transformer has 1,000 turns of wire, and its secondary coil has 2,000 turns of wire. The input voltage is 140 V. What is the output voltage of the transformer? 280 V, 420 V, 60 V , or 70 V

OpenStudy (anonymous):

is it just 140 * 2 =280 V as the solution? :/

OpenStudy (anonymous):

2:1 step up transformer : T = Vs / Vp = 2 ; ⇒ Vs = 2 Vp = 2 × 140 = 280 V

OpenStudy (anonymous):

=280 V

OpenStudy (michele_laino):

we have the subsequent equation: \[\large \frac{{{V_1}}}{{{V_2}}} = \frac{{{N_1}}}{{{N_2}}}\]

OpenStudy (anonymous):

ooh yay thanks!:)

OpenStudy (anonymous):

np